Suppose a random variable is normally distributed. The probabilities for 85 and 142 are 10% and 65%, respectively. Find the mean and standard deviation tp the nearest whole number.

1
Expert's answer
2022-01-04T11:18:03-0500

Solution:

Let the mean and standard deviation be "\,u, \sigma" respectively.

"X\sim N(\mu,\sigma)"


"P(X<85)=P(Z<\dfrac{85-\mu}{\sigma})=0.1""\dfrac{85-\mu}{\sigma}\approx-1.2815516""P(X<142)=P(Z<\dfrac{142-\mu}{\sigma})=0.65""\dfrac{142-\mu}{\sigma}\approx0.385320""\dfrac{85-\mu}{-1.2815516}=\dfrac{142-\mu}{0.385320}""\mu=\dfrac{85(0.385320)+142(1.2815516)}{0.385320+1.2815516}""\mu\approx128.8237""\sigma=\dfrac{142-128.8237}{0.385320}""\sigma=34.1957"

Answer:

"\mu=129, \sigma=34"